OpenStudy (dumbsearch2):

Chris spends $3 per square foot on carpet for a square room. If he spends about $430 on carpet, what is the approximate length, in feet, of a side of the room?

OpenStudy (yanasidlinskiy):

Is this a multiple choice question?

OpenStudy (imstuck):

If he spends $3 per square foot and he spends a total of $430, you can use that info to find out how many square feet total the room is. How do you do that?

OpenStudy (yanasidlinskiy):

Area = L * L = L^2 3 * L^2 = 430 L^2 = 143 so L is about 12 because 12*12 = 144

OpenStudy (yanasidlinskiy):

I'm sure this is how you'd do it!:) I'm not 100% sure. Someone else can see if I've done it right.

OpenStudy (imstuck):

Yes, you did it right. I would have divided the $430 by the 3 to get 143, which is close to 144, and since the room is square, the 12x12 room applies to that just perfectly!

OpenStudy (yanasidlinskiy):

Oh, well there's many ways to figure this out. Soo I guess you're going to come out to the same answer.